OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions -7 and -1

OpenStudy (anonymous):

(x+7)(x+1)=0

jimthompson5910 (jim_thompson5910):

x=-7 or x=-1 x+7=0 or x+1=0 (x+7)(x+1)=0 x^2+1x+7x+7=0 x^2+8x+7=0

OpenStudy (anonymous):

x^2+8x+7=0

OpenStudy (anonymous):

THANK YOU SO MUCH!!!